% Generated by IEEEtran.bst, version: 1.14 (2015/08/26)
\begin{thebibliography}{10}
\providecommand{\url}[1]{#1}
\csname url@samestyle\endcsname
\providecommand{\newblock}{\relax}
\providecommand{\bibinfo}[2]{#2}
\providecommand{\BIBentrySTDinterwordspacing}{\spaceskip=0pt\relax}
\providecommand{\BIBentryALTinterwordstretchfactor}{4}
\providecommand{\BIBentryALTinterwordspacing}{\spaceskip=\fontdimen2\font plus
\BIBentryALTinterwordstretchfactor\fontdimen3\font minus
  \fontdimen4\font\relax}
\providecommand{\BIBforeignlanguage}[2]{{%
\expandafter\ifx\csname l@#1\endcsname\relax
\typeout{** WARNING: IEEEtran.bst: No hyphenation pattern has been}%
\typeout{** loaded for the language `#1'. Using the pattern for}%
\typeout{** the default language instead.}%
\else
\language=\csname l@#1\endcsname
\fi
#2}}
\providecommand{\BIBdecl}{\relax}
\BIBdecl

\bibitem{9140329}
M.~Di~Renzo, A.~Zappone, M.~Debbah, M.-S. Alouini, C.~Yuen, J.~de~Rosny, and
  S.~Tretyakov, ``Smart radio environments empowered by reconfigurable
  intelligent surfaces: How it works, state of research, and the road ahead,''
  \emph{IEEE Journal on Selected Areas in Communications}, vol.~38, no.~11, pp.
  2450--2525, 2020.

\bibitem{9864116}
M.~Di~Renzo, F.~H. Danufane, and S.~Tretyakov, ``Communication models for
  reconfigurable intelligent surfaces: From surface electromagnetics to
  wireless networks optimization,'' \emph{Proceedings of the IEEE}, vol. 110,
  no.~9, pp. 1164--1209, 2022.

\bibitem{9713744}
V.~Degli-Esposti, E.~M. Vitucci, M.~Di~Renzo, and S.~A. Tretyakov,
  ``Reradiation and scattering from a reconfigurable intelligent surface: A
  general macroscopic model,'' \emph{IEEE Transactions on Antennas and
  Propagation}, vol.~70, no.~10, pp. 8691--8706, 2022.

\bibitem{10415512}
M.~Di~Renzo and M.~D. Migliore, ``Electromagnetic signal and information
  theory,'' \emph{IEEE BITS the Information Theory Magazine}, pp. 1--13, 2024.

\bibitem{5446312}
M.~T. Ivrlač and J.~A. Nossek, ``Toward a circuit theory of communication,''
  \emph{IEEE Transactions on Circuits and Systems {I}: Regular Papers},
  vol.~57, no.~7, pp. 1663--1683, 2010.

\bibitem{9319694}
G.~Gradoni and M.~Di~Renzo, ``End-to-end mutual coupling aware communication
  model for reconfigurable intelligent surfaces: An electromagnetic-compliant
  approach based on mutual impedances,'' \emph{IEEE Wireless Communications
  Letters}, vol.~10, no.~5, pp. 938--942, 2021.

\bibitem{10574199}
M.~Nerini, S.~Shen, H.~Li, M.~Di~Renzo, and B.~Clerckx, ``A universal framework
  for multiport network analysis of reconfigurable intelligent surfaces,''
  \emph{IEEE Transactions on Wireless Communications}, vol.~23, no.~10, pp.
  14\,575--14\,590, 2024.

\bibitem{10133771}
M.~Di~Renzo, V.~Galdi, and G.~Castaldi, ``Modeling the mutual coupling of
  reconfigurable metasurfaces,'' in \emph{2023 17th European Conference on
  Antennas and Propagation (EuCAP)}, 2023, pp. 1--4.

\bibitem{10453467}
A.~Rabault, L.~Le~Magoarou, J.~Sol, G.~C. Alexandropoulos, N.~Shlezinger,
  H.~Vincent~Poor, and P.~del Hougne, ``On the tacit linearity assumption in
  common cascaded models of {RIS}-parametrized wireless channels,'' \emph{IEEE
  Transactions on Wireless Communications}, vol.~23, no.~8, pp.
  10\,001--10\,014, 2024.

\bibitem{10195937}
P.~Mursia, S.~Phang, V.~Sciancalepore, G.~Gradoni, and M.~Di~Renzo, ``{SARIS}:
  Scattering aware reconfigurable intelligent surface model and optimization
  for complex propagation channels,'' \emph{IEEE Wireless Communications
  Letters}, vol.~12, no.~11, pp. 1921--1925, 2023.

\bibitem{10476858}
M.~Akrout, F.~Bellili, A.~Mezghani, and J.~A. Nossek, ``Physically consistent
  models for intelligent reflective surface-assisted communications under
  mutual coupling and element size constraint,'' in \emph{2023 57th Asilomar
  Conference on Signals, Systems, and Computers}, 2023, pp. 1589--1594.

\bibitem{9514409}
S.~Shen, B.~Clerckx, and R.~Murch, ``Modeling and architecture design of
  reconfigurable intelligent surfaces using scattering parameter network
  analysis,'' \emph{IEEE Transactions on Wireless Communications}, vol.~21,
  no.~2, pp. 1229--1243, 2022.

\bibitem{10551389}
J.~A. Nossek, D.~Semmler, M.~Joham, and W.~Utschick, ``Physically consistent
  modeling of wireless links with reconfigurable intelligent surfaces using
  multiport network analysis,'' \emph{IEEE Wireless Communications Letters},
  vol.~13, no.~8, pp. 2240--2244, 2024.

\bibitem{10501476}
A.~Abrardo, A.~Toccafondi, and M.~Di~Renzo, ``Analysis and optimization of
  reconfigurable intelligent surfaces based on {S}-parameters multiport network
  theory,'' in \emph{2024 18th European Conference on Antennas and Propagation
  (EuCAP)}, 2024, pp. 1--4.

\bibitem{10418928}
H.~Li, S.~Shen, M.~Nerini, M.~Di~Renzo, and B.~Clerckx, ``Beyond diagonal
  reconfigurable intelligent surfaces with mutual coupling: Modeling and
  optimization,'' \emph{IEEE Communications Letters}, vol.~28, no.~4, pp.
  937--941, 2024.

\bibitem{delhougne2024physicscompliantdiagonalrepresentationwireless}
\BIBentryALTinterwordspacing
P.~del Hougne, ``A physics-compliant diagonal representation for wireless
  channels parametrized by beyond-diagonal reconfigurable intelligent
  surfaces,'' 2024. [Online]. Available: \url{https://arxiv.org/abs/2409.20509}
\BIBentrySTDinterwordspacing

\bibitem{semmler2024performanceanalysissystemscoupled}
\BIBentryALTinterwordspacing
D.~Semmler, J.~A. Nossek, M.~Joham, and W.~Utschick, ``Performance analysis of
  systems with coupled and decoupled {RISs},'' 2024. [Online]. Available:
  \url{https://arxiv.org/abs/2402.15423}
\BIBentrySTDinterwordspacing

\bibitem{10511924}
J.~A. Nossek, D.~Semmler, M.~Joham, and W.~Utschick, ``Modelling of wireless
  links with reconfigurable intelligent surfaces using multiport network
  analysis,'' in \emph{2024 27th International Workshop on Smart Antennas
  (WSA)}, 2024, pp. 1--8.

\bibitem{10501453}
S.~I. Raptis, A.~Papadopoulos, L.~Symeonidis, A.~Lalas, C.~K. Liaskos,
  K.~Votis, D.~Tzovaras, and T.~V. Yioultsis, ``An accurate semi-analytical
  model for periodic tunable metasurfaces electromagnetic response,'' in
  \emph{2024 18th European Conference on Antennas and Propagation (EuCAP)},
  2024, pp. 1--5.

\bibitem{10008536}
R.~Faqiri, C.~Saigre-Tardif, G.~C. Alexandropoulos, N.~Shlezinger, M.~F. Imani,
  and P.~D. Hougne, ``End-to-end fading channel modeling for {RIS}-empowered
  smart wireless environments,'' in \emph{2022 IEEE Globecom Workshops (GC
  Wkshps)}, 2022, pp. 741--746.

\bibitem{9856592}
R.~Faqiri, C.~Saigre-Tardif, G.~C. Alexandropoulos, N.~Shlezinger, M.~F. Imani,
  and P.~del Hougne, ``Physfad: Physics-based end-to-end channel modeling of
  {RIS}-parametrized environments with adjustable fading,'' \emph{IEEE
  Transactions on Wireless Communications}, vol.~22, no.~1, pp. 580--595, 2023.

\bibitem{10501073}
P.~Del~Hougne, ``Load impedances vs polarizabilities: On the compactness of
  physics-compliant models of {RIS}-parametrized wireless channels,'' in
  \emph{2024 18th European Conference on Antennas and Propagation (EuCAP)},
  2024, pp. 1--4.

\bibitem{10310003}
H.~Prod’homme and P.~del Hougne, ``Efficient computation of physics-compliant
  channel realizations for (rich-scattering) {RIS}-parametrized radio
  environments,'' \emph{IEEE Communications Letters}, vol.~27, no.~12, pp.
  3375--3379, 2023.

\bibitem{delhougne2023risparametrizedrichscatteringenvironmentsphysicscompliant}
\BIBentryALTinterwordspacing
P.~del Hougne, ``{RIS}-parametrized rich-scattering environments:
  Physics-compliant models, channel estimation, and optimization,'' 2023.
  [Online]. Available: \url{https://arxiv.org/abs/2311.11651}
\BIBentrySTDinterwordspacing

\bibitem{9360851}
X.~Qian and M.~Di~Renzo, ``Mutual coupling and unit cell aware optimization for
  reconfigurable intelligent surfaces,'' \emph{IEEE Wireless Communications
  Letters}, vol.~10, no.~6, pp. 1183--1187, 2021.

\bibitem{9525465}
A.~Abrardo, D.~Dardari, M.~Di~Renzo, and X.~Qian, ``{MIMO} interference
  channels assisted by reconfigurable intelligent surfaces: Mutual coupling
  aware sum-rate optimization based on a mutual impedance channel model,''
  \emph{IEEE Wireless Communications Letters}, vol.~10, no.~12, pp. 2624--2628,
  2021.

\bibitem{10666074}
A.~Abrardo, A.~Toccafondi, and M.~Di~Renzo, ``Design of reconfigurable
  intelligent surfaces by using {S}-parameter multiport network theory --
  {O}ptimization and full-wave validation,'' \emph{IEEE Transactions on
  Wireless Communications}, pp. 1--1, 2024.

\bibitem{wijekoon2024electromagneticallyconsistentmodelingoptimizationmutual}
\BIBentryALTinterwordspacing
D.~Wijekoon, A.~Mezghani, G.~C. Alexandropoulos, and E.~Hossain,
  ``Electromagnetically-consistent modeling and optimization of mutual coupling
  in {RIS}-assisted multi-user {MIMO} communication systems,'' 2024. [Online].
  Available: \url{https://arxiv.org/abs/2404.04539}
\BIBentrySTDinterwordspacing

\bibitem{10666709}
A.~Abrardo, ``Optimizing reconfigurable intelligent surfaces in multi-user
  environments: A multiport network theory approach leveraging statistical
  {CSI},'' \emph{IEEE Transactions on Communications}, pp. 1--1, 2024.

\bibitem{10476915}
N.~S. Perović, L.-N. Tran, M.~Di~Renzo, and M.~F. Flanagan, ``Optimization of
  {RIS}-aided {SISO} systems based on a mutually coupled loaded wire dipole
  model,'' in \emph{2023 57th Asilomar Conference on Signals, Systems, and
  Computers}, 2023, pp. 145--150.

\bibitem{10352994}
H.~E. Hassani, X.~Qian, S.~Jeong, N.~S. Perović, M.~Di~Renzo, P.~Mursia,
  V.~Sciancalepore, and X.~Costa-Pérez, ``Optimization of {RIS}-aided {MIMO}
  -- a mutually coupled loaded wire dipole model,'' \emph{IEEE Wireless
  Communications Letters}, vol.~13, no.~3, pp. 726--730, 2024.

\bibitem{10576064}
K.~Konno, S.~Terranova, Q.~Chen, and G.~Gradoni, ``Generalized impedance model
  of wireless links assisted by reconfigurable intelligent surfaces,''
  \emph{IEEE Transactions on Antennas and Propagation}, vol.~72, no.~10, pp.
  7691--7699, 2024.

\bibitem{10246382}
G.~Pettanice, R.~Valentini, P.~D. Marco, F.~Loreto, D.~Romano, F.~Santucci,
  D.~Spina, and G.~Antonini, ``Mutual coupling aware time-domain
  characterization and performance analysis of reconfigurable intelligent
  surfaces,'' \emph{IEEE Transactions on Electromagnetic Compatibility},
  vol.~65, no.~6, pp. 1606--1620, 2023.

\bibitem{9901291}
G.~Pettanice, F.~Loreto, P.~Di~Marco, D.~Romano, F.~Santucci, R.~Alesii, and
  G.~Antonini, ``Time-domain characterization of reconfigurable intelligent
  surfaces for wireless communications,'' in \emph{2022 International Symposium
  on Electromagnetic Compatibility – EMC Europe}, 2022, pp. 566--571.

\bibitem{pettanice2024designoptimizationreconfigurableintelligent}
\BIBentryALTinterwordspacing
G.~Pettanice, M.~{Di Renzo}, R.~Valentini, S.~Jeong, P.~{Di Marco},
  F.~Santucci, D.~Romano, and G.~Antonini, ``Design and optimization of
  reconfigurable intelligent surfaces using the {PEEC} method,'' 2024.
  [Online]. Available: \url{https://arxiv.org/abs/2404.18315}
\BIBentrySTDinterwordspacing

\bibitem{pettanice2024multiportnetworkmodelingreconfigurable}
\BIBentryALTinterwordspacing
G.~Pettanice, M.~{Di Renzo}, S.~Jeong, R.~Valentini, P.~{Di Marco},
  F.~Santucci, D.~Romano, and G.~Antonini, ``Multiport network modeling for
  reconfigurable intelligent surfaces: Numerical validation with a full-wave
  {PEEC} simulator,'' 2024. [Online]. Available:
  \url{https://arxiv.org/abs/2404.18310}
\BIBentrySTDinterwordspacing

\bibitem{10528316}
J.~Tapie, H.~Prod’homme, M.~F. Imani, and P.~d. Hougne, ``Systematic
  physics-compliant analysis of over-the-air channel equalization in
  {RIS}-parametrized wireless networks-on-chip,'' \emph{IEEE Journal on
  Selected Areas in Communications}, vol.~42, no.~8, pp. 2026--2038, 2024.

\bibitem{sol2023experimentallyrealizedphysicalmodelbasedwave}
J.~Sol, H.~Prod'homme, L.~Le~Magoarou, and P.~Hougne, ``Experimentally realized
  physical-model-based frugal wave control in metasurface-programmable complex
  media,'' 04 2024.

\bibitem{10501200}
P.~Mursia, T.~Mazloum, F.~Munoz, V.~Sciancalepore, G.~Gradoni, R.~D'Errico,
  M.~Di~Renzo, X.~Costa-Pérez, A.~Clemente, and G.~Lerosey, ``Empirical
  validation of the impedance-based {RIS} channel model in an indoor scattering
  environment,'' in \emph{2024 18th European Conference on Antennas and
  Propagation (EuCAP)}, 2024, pp. 1--4.

\bibitem{10669140}
P.~Zheng, R.~Wang, A.~Shamim, and T.~Y. Al-Naffouri, ``Mutual coupling in
  {RIS}-aided communication: Model training and experimental validation,''
  \emph{IEEE Transactions on Wireless Communications}, pp. 1--1, 2024.

\bibitem{delhougne2024virtualvnaminimalambiguityscattering}
\BIBentryALTinterwordspacing
P.~del Hougne, ``Virtual {VNA}: Minimal-ambiguity scattering matrix estimation
  with load-tunable ports,'' 2024. [Online]. Available:
  \url{https://arxiv.org/abs/2403.08074}
\BIBentrySTDinterwordspacing

\bibitem{delhougne2024virtualvna20ambiguityfree}
\BIBentryALTinterwordspacing
------, ``Virtual {VNA} 2.0: Ambiguity-free scattering matrix estimation by
  terminating inaccessible ports with tunable and coupled loads,'' 2024.
  [Online]. Available: \url{https://arxiv.org/abs/2409.10977}
\BIBentrySTDinterwordspacing

\bibitem{9818951}
Z.~Zhang, J.~W. Zhang, J.~W. Wu, J.~C. Liang, Z.~X. Wang, Q.~Cheng, Q.~S.
  Cheng, T.~J. Cui, H.~Q. Yang, G.~B. Liu, and S.~R. Wang, ``Macromodeling of
  reconfigurable intelligent surface based on microwave network theory,''
  \emph{IEEE Transactions on Antennas and Propagation}, vol.~70, no.~10, pp.
  8707--8717, 2022.

\bibitem{10440504}
P.~Zheng, X.~Ma, and T.~Y. Al-Naffouri, ``On the impact of mutual coupling on
  {RIS}-assisted channel estimation,'' \emph{IEEE Wireless Communications
  Letters}, vol.~13, no.~5, pp. 1275--1279, 2024.

\bibitem{arXiv.2410.04110}
P.~Zheng, S.~Tarboush, H.~Sarieddeen, and T.~Al-Naffouri, ``Mutual
  coupling-aware channel estimation and beamforming for {RIS}-assisted
  communications,'' 10 2024.

\bibitem{10279095}
R.~Ma, J.~Tang, X.~Zhang, K.-K. Wong, and J.~A. Chambers, ``Joint active
  beamforming and circuit parameter optimization for reconfigurable intelligent
  surface-aided {SWIPT} systems,'' in \emph{ICC 2023 - IEEE International
  Conference on Communications}, 2023, pp. 937--942.

\bibitem{10032536}
------, ``Energy-efficiency optimization for mutual-coupling-aware wireless
  communication system based on {RIS}-enhanced {SWIPT},'' \emph{IEEE Internet
  of Things Journal}, vol.~10, no.~22, pp. 19\,399--19\,414, 2023.

\bibitem{10375247}
R.~Ma, J.~Tang, X.~Y. Zhang, K.-K. Wong, and J.~A. Chambers, ``{RIS}-assisted
  {SWIPT} network for internet of everything under the electromagnetics-based
  communication model,'' \emph{IEEE Internet of Things Journal}, vol.~11,
  no.~9, pp. 15\,402--15\,415, 2024.

\bibitem{10606051}
R.~Ma, J.~Tang, X.~Yin~Zhang, K.-K. Wong, and J.~A. Chambers, ``{RIS}-assisted
  hybrid symbiotic radio network: From practical end-to-end model to
  multi-objective optimization,'' \emph{IEEE Transactions on Cognitive
  Communications and Networking}, vol.~10, no.~5, pp. 1730--1743, 2024.

\bibitem{peng2024risnetdomainknowledgedrivenneural}
\BIBentryALTinterwordspacing
B.~Peng, K.-L. Besser, S.~Shen, F.~Siegismund-Poschmann, R.~Raghunath,
  D.~Mittleman, V.~Jamali, and E.~A. Jorswieck, ``{RISnet}: A domain-knowledge
  driven neural network architecture for {RIS} optimization with mutual
  coupling and partial {CSI},'' 2024. [Online]. Available:
  \url{https://arxiv.org/abs/2403.04028}
\BIBentrySTDinterwordspacing

\bibitem{10530995}
M.~Nerini and B.~Clerckx, ``Physically consistent modeling of stacked
  intelligent metasurfaces implemented with beyond diagonal {RIS},'' \emph{IEEE
  Communications Letters}, vol.~28, no.~7, pp. 1693--1697, 2024.

\bibitem{10636283}
P.~Mursia, F.~Devoti, M.~Rossanese, V.~Sciancalepore, G.~Gradoni, M.~D. Renzo,
  and X.~Costa-Pérez, ``{T3DRIS}: Advancing conformal {RIS} design through
  in-depth analysis of mutual coupling effects,'' \emph{IEEE Transactions on
  Communications}, pp. 1--1, 2024.

\bibitem{10515204}
J.~An, C.~Yuen, C.~Xu, H.~Li, D.~W.~K. Ng, M.~Di~Renzo, M.~Debbah, and
  L.~Hanzo, ``Stacked intelligent metasurface-aided {MIMO} transceiver
  design,'' \emph{IEEE Wireless Communications}, vol.~31, no.~4, pp. 123--131,
  2024.

\end{thebibliography}

\documentclass[conference,a4paper]{IEEEtran}
\usepackage{xcolor}
\usepackage{balance}

\ifCLASSINFOpdf
  \usepackage[pdftex]{graphicx}
\else
  \usepackage[dvips]{graphicx}
\fi

\ifCLASSOPTIONcompsoc
 \usepackage[caption=false,font=normalsize,labelfont=sf,textfont=sf]{subfig}
\else
 \usepackage[caption=false,font=footnotesize]{subfig}
\fi

\begin{document}

\title{Multiport Network Theory for Modeling and Optimizing Reconfigurable Metasurfaces}

\author{\IEEEauthorblockN{Marco Di Renzo\IEEEauthorrefmark{1} and   
Philipp del Hougne\IEEEauthorrefmark{2}   
}                                     
\IEEEauthorblockA{\IEEEauthorrefmark{1}Universit\'e Paris-Saclay, CNRS, CentraleSup\'elec \\ Laboratoire des Signaux et Syst\'emes, 91192 Gif-sur-Yvette, France}
\IEEEauthorblockA{\IEEEauthorrefmark{2}Universit\'e de Rennes, CNRS, IETR -- UMR 6164, F-35000
Rennes, France}
\IEEEauthorblockA{ \emph{marco.di-renzo@universite-paris-saclay.fr} }
}

\maketitle

\begin{abstract}
Multiport network theory (MNT) is a powerful analytical tool for modeling and optimizing complex systems based on circuit models. We present an overview of current research on the application of MNT to the development of electromagnetically consistent models for programmable metasurfaces, with focus on reconfigurable intelligent surfaces for wireless communications.
\end{abstract}

\vskip0.5\baselineskip
\begin{IEEEkeywords}
Smart radio environments, reconfigurable intelligent surfaces, multiport network theory, circuits theory.
\end{IEEEkeywords}

\section{Introduction}
Reconfigurable intelligent surface (RIS) is an emerging technology in wireless communications, which turns wireless networks into programmable environments thanks to the wave manipulation capabilities of programmable metasurfaces \cite{9140329}. One of the main open research challenges in wireless communication theory is the development of electromagnetically consistent models for analyzing and optimizing RIS-aided communication networks \cite{9864116}, \cite{9713744}. This need has recently motivated communication theorists to consider the reintroduction of electromagnetics in communication theory, with special focus on surface electromagnetics when RIS is considered \cite{10415512}.

In this context, circuits theory is viewed as a powerful analytical tool for modeling complex communication systems, as it offers a formal theory and efficient methods for analysis and optimization \cite{5446312}. More specifically, multiport network theory (MNT) constitutes a convenient abstraction model as the perfomance metrics of interest can be expressed in a matrix form, which resembles well-known models utilized in signal processing for multiple-input multiple-antenna (MIMO) systems. The general idea of MNT is that each network element, whether it is a transmitter, a receiver, or a scattering object, can be modeled as a multiport circuit whose ports are appropriately loaded. The first application of MNT to RIS-aided networks was reported in  \cite{9319694}. A recent tutorial on MNT applied to RIS-aided networks can be found in \cite{10574199}.

\section{State of the Art}
In this section, we classify available research works on
MNT with focus on RIS-aided wireless communications.

\subsection{Modeling}
As mentioned, the first application of MNT to RIS-aided networks is reported in \cite{9319694}. The proposed approach is based on impedance parameters. A closed-form expression for the mutual impedances is given in \cite{10133771}, by assuming that the RIS elements can be approximated as loaded dipoles. As elaborated in \cite{10453467}, a model based on MNT leads to a non-linear end-to-end response as a function of the optimization variables. In \cite{10195937}, the MNT model is generalized for application to multipath channels, by representing general scattering objects as multiport networks. In \cite{10476858}, the authors extend the work in \cite{9319694}, by considering Chu's theory. The authors of \cite{9514409}, \cite{10551389} and \cite{10501476} propose the use of scattering parameters in lieu of the impedance parameters, and the structural scattering of RIS is discussed in the latter two papers. The application to fully-connected (non-diagonal) impedance and scattering parameters is analyzed in \cite{10418928}, and it is shown that these architectures are beneficial in the presence of electromagnetic mutual coupling. In \cite{delhougne2024physicscompliantdiagonalrepresentationwireless},  the author introduces a physics-compliant diagonal representation for RISs that are parametrized with non-diagonal impedance matrices. In \cite{semmler2024performanceanalysissystemscoupled}, the authors introduce the concept of decoupling network. In \cite{10511924}, the benefits of utilizing the impedance parameters from the physics point of view are discussed. An MNT model for periodic RISs that account for Floquet's modes is introduced in \cite{10501453}. Finally, the authors of \cite{10008536},
\cite{9856592}, \cite{10501073}, \cite{10310003} have introduced a model based on the coupled dipoles formalism, which is expressed in terms of polarizabilities. A comparison between MNT and the coupled dipoles formalism is available  in \cite{delhougne2023risparametrizedrichscatteringenvironmentsphysicscompliant}.

\subsection{Optimization}
Several optimization algorithms based on MNT models have been proposed recently. The first optimization algorithm was reported in \cite{9360851} for single-antenna transmitters and receivers. The approach is based on the Neumann series approximation in order to efficiently handle the inherent non-linearly of the model. The approach was then utilized in several other papers \cite{10195937}, \cite{9525465}, \cite{10666074}, \cite{wijekoon2024electromagneticallyconsistentmodelingoptimizationmutual}, \cite{10666709}. In \cite{10666709}, in particular, the authors propose the first approach to control the structural scattering by design, and they show that the representation in terms of scattering parameters leads to a faster convergence rate. Optimization methods that avoid the Neumann series approximation are reported in \cite{10476915} and \cite{10352994}. The approach in \cite{10352994}, in particular, can be applied to general multiple-antenna systems.

\subsection{Validation Using Full-Wave Simulations}
Given the flexibility of MNT for modeling and optimizing RIS-aided wireless systems, and, in particular, near-field aspects, such as the electromagnetic mutual coupling, a few research works have recently analyzed the accuracy of communication models obtained by using MNT with the aid of electromagnetic simulators \cite{10666074}, \cite{10576064}, \cite{10246382}, \cite{9901291},  \cite{pettanice2024designoptimizationreconfigurableintelligent},
\cite{pettanice2024multiportnetworkmodelingreconfigurable}, \cite{10528316}. In \cite{10666074}, in particular, MNT is shown to be able to model the structural scattering in a consistent manner. In \cite{pettanice2024designoptimizationreconfigurableintelligent} and
\cite{pettanice2024multiportnetworkmodelingreconfigurable}, in addition, an efficient electromagnetic simulator based on the partial element equivalent circuit method is proposed. In \cite{10528316}, it was shown that a compact MNT-model can be calibrated with a single full-wave simulation even in rich-scattering environments involving metallic walls, extended dielectric objects and significant RIS element structural scattering, highlighting that these scattering entities do not have to be explicitly represented by additional ports, if this is not required for design and optimization purposes.

\subsection{Validation Using Measurements and Testbeds}

Besides electromagnetic simulations, there have been experimental efforts to validate physics-compliant models for RIS-parametrized channels. This oftentimes amounts to end-to-end physics-compliant channel estimation. The first report on this problem \cite{sol2023experimentallyrealizedphysicalmodelbasedwave} was in fact not based on MNT but on an operationally equivalent coupled-dipole formalism. The considered radio environment involves unknown rich scattering, but the results in \cite{sol2023experimentallyrealizedphysicalmodelbasedwave} highlight that there is no need to explicitly model the rich scattering as long as its influence on the coupling between the primary entities (antennas and RIS elements) is correctly estimated. Thereby, the number of parameters to be estimated is kept minimal. In \cite{sol2023experimentallyrealizedphysicalmodelbasedwave}, a gradient-descent approach is utilized to estimate the model parameters and it is further highlighted that this is possible based on non-coherent detection, which greatly alleviates the measurement burden. Efforts involving MNT models are reported in \cite{10501200}, \cite{10669140}, \cite{delhougne2024virtualvnaminimalambiguityscattering}, \cite{delhougne2024virtualvna20ambiguityfree}, \cite{9818951}. In \cite{10501200}, the scattering objects in a multipath environment are explicitly modeled in a statistical manner. In a quest to minimize the number of model parameters and their ambiguities, the authors of \cite{delhougne2024virtualvnaminimalambiguityscattering} and  \cite{delhougne2024virtualvna20ambiguityfree} demonstrated experimentally that at least three distinct and known states per RIS element and coupling between neighboring RIS elements are required, reporting closed-form and gradient-descent based approaches. The same methods were shown to be directly applicable to fully-connected RIS in an experimentally grounded case-study in \cite{delhougne2024physicscompliantdiagonalrepresentationwireless}, by leveraging a physics-compliant MNT-based diagonal representation of fully-connected RISs.  Meanwhile, \cite{10669140} and \cite{9818951} considered a free-space radio environment and demonstrated that parameters estimated in simulation are validated experimentally.

\subsection{Applications in Wireless Systems}
Due to the advantageous tradeoff in terms of electromagnetic accuracy and suitability for optimization, MNT has been utilized for various communication problems. In \cite{10440504} and \cite{arXiv.2410.04110}, it is utilized to evaluate the impact of mutual coupling for channel estimation. In \cite{10279095}, \cite{10032536}, \cite{10375247}, it is utilized to analyze the energy efficiency of wireless systems with simultaneous information and power transfer. Other applications include the analysis and optimization of symbiotic radio systems \cite{10606051}, wireless networks on chip \cite{10528316}, machine learning optimization of scattering parameters \cite{peng2024risnetdomainknowledgedrivenneural}, modeling multi-layer transmissive metasurfaces \cite{10530995}, and the optimization of conformal metasurfaces \cite{10636283}.

\balance

\section{Conclusion}
In this paper, we have summarized the state of the art on the application of MNT for modeling and optimizing RIS-aided wireless communications. Current results have highlighted that MNT is a powerful tool for modeling and optimizing smart radio environments, whose accuracy has been validated with the aid of electromagnetic simulations and empirical measurements. For analytical tractability, most works consider loaded dipoles to model the RIS elements. On the other hand, more complex structures for the unit cells are handled by using semi-analytical methods. In light of the promising tradeoff in terms of analytical tractability, suitability for optimization, accuracy, and engineering insights, more efforts are needed to develop MNT-based models that do not rely on electromagnetic simulations and that can be utilized for general reconfigurable metasurfaces integrating communication, sensing and computing in the wave domain \cite{10515204}.

\section*{Acknowledgment}
The work of M. Di Renzo was supported in part by the European Commission through the Horizon Europe project COVER under grant agreement 101086228, the Horizon Europe project UNITE under grant agreement 101129618, and the Horizon Europe project INSTINCT under grant agreement 101139161, as well as by the Agence Nationale de la Recherche (ANR) through the France 2030 project ANR-PEPR Networks of the Future under grant agreements NF-PERSEUS 22-PEFT-0004, NF-YACARI 22-PEFT-0005, NF-SYSTERA 22-PEFT-0006, NF-FOUND 22-PEFT-0010, and by the CHIST-ERA project PASSIONATE under grant agreements CHIST-ERA-22-WAI-04 and ANR-23-CHR4-0003-01. The work of P. del Hougne was supported in part by the Agence Nationale de la Recherche (ANR) through the France 2030 project ANR-PEPR Networks of the Future under grant agreement NF-YACARI 22-PEFT-0005.

\bibliographystyle{IEEEtran}
\bibliography{biblio}

\end{document}